OpenStudy (anonymous):

Which of the following is equivalent to log125 25 ?

OpenStudy (anonymous):

\[\log _{125}25\]

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

1.076 3/2 2/3 2/5

jimthompson5910 (jim_thompson5910):

\[\log_{125}(25)\] \[\frac{\log_{10}(25)}{\log_{10}(125)}\] \[\frac{\log_{10}(5^2)}{\log_{10}(5^3)}\] \[\frac{2\log_{10}(5)}{3\log_{10}(5)}\] \[\frac{2}{3}\] So \[\log_{125}(25)=\frac{2}{3}\]